OpenStudy (anonymous):

a hemisphere has a radius of 8.5 inches. what is the radius of a sphere with the same surface area? to the nearest tenth of an inch

OpenStudy (anonymous):

\[hemisphereSA=2 \pi r^2\] \[sphereSA=4 \pi r^2\]

OpenStudy (anonymous):

\[hemisphereSA = 2 \pi 8.5^2--->144.5 \pi\] \[sphereSA=4 \pi r^2=144.5 \pi\] \[r^2=\frac{ 144.5 \pi }{ 4 \pi }\] \[r^2=\frac{ 144.5 }{ 4 }\] \[r=\sqrt{\frac{ 144.5 }{ 4 }}\] \[r=6.01\]